I NEED HELP ON THIS PROBLEM {{ x+y=10}}
{{ y=x-4}}

in the book doesent explain it well and when mt teacher tells me how to do it its confusing! can u help me
thank you!

Answer by stanbon(48568)

1st: x+y=10
2nd: y=x-4
--------
Substitute the value for y into the 1st equation, as follows:
x+(x-4)=10
Simplify to get:
2x-4=10
2x=14
x=7
Now, substitute that value of x into the 2nd equation to solve for y, as follows:
y=7-4=3
---------
Now check those answers in the 1st equation as follows:
7+3=10
10=10
----------
Solution:
x=7 and y=3 is the solution for the system of equations.
Cheers,
Stan H.